mays is a diploid species with 10 pairs of chromosomes in each somatic cell. How many chromosomes will be found per cell in each of the following chromosome mutants? Questions 64 to 67 are connected 64) Trisomy? E) 21 D) 30 A) 19 B) 22 65) Monosomy E) 21 A) 19 B) 22 C) 60 66) triploid E) 21 B) 22 67) Tetrasomy D) 30 E) 21 A) 19 B) 22 68) Species I has 2n -12 chromosomes and species I has 2n-14 chromosomes. What would be the expected chromosome numbers in individual organisms with the following chromosome mutations? Allotriploidy including species I and II A) 12 or 14 D) 6 or 7 ?) 20 and 30 B)24 or 28 d) 190r 20 59) These scientists conducted experiment with N14 and N15 isotopes and discovered that DNA replica is semi-conservative A) Watson and Crick D) Watson and Chargaff B) Meselson and Stahl C) Avery and McCathy E) Rosalind Franklin and Maurice Wilkins 70) This scientist injected mice with the two strains of bacteria, one virulent (S) and the other aviruler (R). The S bacteria strain killed mice whereas the avirulent R bacteria did not. When S strain was he killed and mixed with R strain and injected to mice, it died. Thus the scientist discovered transformat A) Chargaff B) Rosalin C) Watson D) Griffith E) Crick 71) These evidence was available to Watson and Crick that helped them to discover the struc of DNA A) Rosalind Franklin X-ray diffraction studies B) Chargaff determination that the amount of A equals T and G equals C C) Linus Pauling deduce the structure of an alpha helix in protein D) all of the above E) none of the above 72) A virus contains 10% adenine, 1090 uracil, 41% guanine, and 39% cytosine. Double-stranded DNA B) Single-stranded DNA C) Double-stranded RNA, o types of amino acids are found in the amino terminal region of histone mol C) serine and tyrosine This virus is: D) Single-stranded RNA E) Chloroplast DNA 3) These tw A) L eucine and Isoleucine B) glycine and lysine tophan and glycine E) arginine and lysine ec
Answer #1

Zea mays:

2n = 20

n = 10

64. Trisomy = One extra copy of a chromosome = 2n + 1 = 21

65. Monosomy = Absence of a chromosome = 2n - 1 = 19

66. Triploid = 3n = 30

67. Tetrasomy = 2n + 2 = 22

68. Allotriploid = n from one species and 2n from another species = 6 + 14 or 12 + 7 = 19 or 20

69. Messelson and Stahl conducted experiments with radioisotopes to confirm the semiconservative nature of DNA replication.

70. Griffith performed transformation experiments on Streptococcus.

71. Watson and Crick used the data from Franklin, Chargaff, and Pauling.

72. The virus contains Uracil and does not follow Chargaff's rule. So, it must be a ssRNA virus.

73. Lysine and ARginine are found at the Histone tails.